OpenStudy (anonymous):

Help please...Use an Addition or Subtraction Formula to write the expression as a trigonometric function of one number. cos 12° cos 18° − sin 12° sin 18°

OpenStudy (anonymous):

\[\cos a \cos b-\sin a \sin b=\cos \left( a+b \right)\]